# MusicMagus: Zero-Shot Text-to-Music Editing via Diffusion Models

### v1

@@ Abstract
Recent advances in text-to-music generation models have opened new avenues in musical creativity. However, music generation usually involves iterative refinements, and how to edit the generated music remains a significant challenge. This paper introduces a novel approach to the editing of music generated by such models, enabling the modification of specific attributes, such as genre, mood and instrument, while maintaining other aspects unchanged. Our method transforms text editing to latent space manipulation while adding an extra constraint to enforce consistency. It seamlessly integrates with existing pretrained text-to-music diffusion models without requiring additional training. Experimental results demonstrate superior performance over both zero-shot and certain supervised baselines in style and timbre transfer evaluations. Additionally, we showcase the practical applicability of our approach in real-world music editing scenarios.  [^1] Project page: https://bit.ly/musicmagus-demo.  [^2] Work done during Yixiao’s internship at Sony AI.
@@ 2.2 Text-to-music editing
[…]
@@ 3 Background
MusicMagus utilizes a pretrained diffusion model Ho et al. (2020) for text-to-music editing, eliminating the need for additional training. Specifically, we use a pretrained AudioLDM 2 model Liu et al. (2023a) as the backbone model. AudioLDM 2 employs a variational autoencoder (VAE) Kingma and Welling (2013) to compress a music audio spectrogram into a latent low-dimensional space. It then trains a latent diffusion model (LDM) on this latent space to generate new samples from Gaussian noise conditioned on text inputs. During generation, the LDM takes a condition $y$, generates a latent variable $z_{0}$, and uses the VAE decoder to produce the music spectrogram $x$. This spectrogram can then be converted into a waveform using an external vocoder, such as HiFi-GAN Su et al. (2020).
@@
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\epsilon_{\theta}(z_{t},t)$ is a neural network that predicts the noise added at step $t$. The network $\epsilon_{\theta}$ is trained to minimize the difference between the predicted noise and the actual noise added during the forward process.
@@ 5.5.1 Subjective experiments
We conducted a subjective listening test for both the timbre transfer and style transfer tasks. This test involved disseminating an online survey within the Music Information Retrieval (MIR) community and our broader research network, which resulted in the collection of 26 complete responses. The gender distribution of the participants was 19 males (76%) and 6 females (24%). Regarding musical experience, 5 participants (19.23%) had less than 1 year of experience, 5 (19.23%) had between 1 and 5 years, and the majority, 16 participants (61.54%), had more than 5 years of experience. This subjective test was approved by the ethics committee of our University (omitted for anonymous review).

### v2

@@ Abstract
Recent advances in text-to-music generation models have opened new avenues in musical creativity. However, music generation usually involves iterative refinements, and how to edit the generated music remains a significant challenge. This paper introduces a novel approach to the editing of music generated by such models, enabling the modification of specific attributes, such as genre, mood and instrument, while maintaining other aspects unchanged. Our method transforms text editing to latent space manipulation while adding an extra constraint to enforce consistency. It seamlessly integrates with existing pretrained text-to-music diffusion models without requiring additional training. Experimental results demonstrate superior performance over both zero-shot and certain supervised baselines in style and timbre transfer evaluations. We also show the practical applicability of our approach in real-world music editing scenarios.  [^1] Code and demo: https://bit.ly/musicmagus-demo.  [^2] Work done during Yixiao’s internship at Sony AI.
@@ 2.2 Text-to-music editing
Compared to text-based image editing  Hertz et al. (2022); Parmar et al. (2023); Hu et al. (2024), research on text-to-music editing is relatively limited. Models like InstructME Han et al. (2023) and M2UGen Hussain et al. (2023) demonstrate capabilities in both inter-stem and intra-stem editing, allowing for structural changes and detailed modifications within stems, but they often require extra training and specific data. Loop Copilot Zhang et al. (2023), an AI agent, employs a combination of existing models to facilitate compositional editing, yet it does so without altering the fundamental architecture or interface of the original models. Furthermore, the tag-conditioned timbre Wu et al. (2023c) and style Dai et al. (2018) transfer task can be seen as a precursor to this task, however labels are not as flexible as text.
@@ 3 Background
MusicMagus utilizes a pretrained diffusion model Ho et al. (2020) for text-to-music editing, eliminating the need for additional training. Specifically, we use a pretrained AudioLDM 2 model Liu et al. (2023a) as the backbone model, since it is the most accessible text-to-music diffusion model with publicly available weights. AudioLDM 2 employs a variational autoencoder (VAE) Kingma and Welling (2013) to compress a music audio spectrogram into a latent low-dimensional space. It then trains a latent diffusion model (LDM) on this latent space to generate new samples from Gaussian noise conditioned on text inputs. During generation, the LDM takes a condition $y$, generates a latent variable $z_{0}$, and uses the VAE decoder to produce the music spectrogram $x$. This spectrogram can then be converted into a waveform using an external vocoder, such as HiFi-GAN Su et al. (2020).
@@
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\epsilon_{\theta}(z_{t},t)$ is a neural network that predicts the noise added at step $t$, and $sigma_{t}$ represents the standard deviation of the noise added. The network $\epsilon_{\theta}$ is trained to minimize the difference between the predicted noise and the actual noise added during the forward process.
@@ 5.5.1 Subjective experiments
We conducted a subjective listening test for both the timbre transfer and style transfer tasks. This test involved disseminating an online survey within the Music Information Retrieval (MIR) community and our broader research network, which resulted in the collection of 26 complete responses. The gender distribution of the participants was 19 males (76%) and 6 females (24%). Regarding musical experience, 5 participants (19.23%) had less than 1 year of experience, 5 (19.23%) had between 1 and 5 years, and the majority, 16 participants (61.54%), had more than 5 years of experience. This subjective test was approved by the ethics committee of both Sony AI and Queen Mary University of London (QMERC20.565.DSEECS23.129).
@@ 7 Conclusion
Ethical Statement
The subjective tests are approved by the ethics committee of both Sony AI and Queen Mary University of London (QMERC20.565.DSEECS23.129).
Acknowledgments
We want to thank Dr. Woosung Choi, Zineb Lahrichi, Liwei Lin and Liming Kuang for their contribution to this work. Yixiao Zhang is a research student at the UKRI Centre for Doctoral Training in Artificial Intelligence and Music, supported jointly by the China Scholarship Council, Queen Mary University of London and Apple Inc.